OpenStudy (anonymous):

An internet retailer conducts a study to determine the proportion of Australians over the age of 25 who have made on-line purchases of consumer durable goods in the last year. The managing director claims the proportion is 15%. A random sample of size 173 is selected. If the claim is valid, below what value would the lowest 10% of sample proportions lie? Express the answer as a percentage, correct to two decimal places.

OpenStudy (anonymous):

pls answer

OpenStudy (anonymous):

@dumbcow pls help

OpenStudy (anonymous):

@fthrogers help

OpenStudy (anonymous):

@Mimi_x3 help

OpenStudy (anonymous):

@lgbasallote help!

OpenStudy (anonymous):

@zepp help

OpenStudy (zepp):

I don't understand this question, sorry. :s The director claims 15%, 15% of how many people?

OpenStudy (anonymous):

1

OpenStudy (anonymous):

@zepp ^^

OpenStudy (zepp):

1 person? o.o

OpenStudy (dumbcow):

hold on, let me get my stats book

OpenStudy (dumbcow):

what is the confidence level of the hypothesis test?

OpenStudy (dumbcow):

ok i'll assume its 0.1 ...(i don't know what the 10% is referring to) z-score for proportions is \[\large z = \frac{p-0.15}{\sqrt{\frac{.15(1-.15)}{173}}}\] at 90% significance, claim is valid if Z>-1.28 \[\rightarrow \frac{p-.15}{.02715} > -1.28\] \[\rightarrow p > .115\]

OpenStudy (dumbcow):

i am also assuming the hypothesis test is: H0: p = 0.15 Ha: p < 0.15 is that correct?